### v2

# OmniGen: Unified Image Generation

Abstract
The emergence of Large Language Models (LLMs) has unified language generation tasks and revolutionized human-machine interaction. However, in the realm of image generation, a unified model capable of handling various tasks within a single framework remains largely unexplored. In this work, we introduce OmniGen, a new diffusion model for unified image generation. OmniGen is characterized by the following features: 1) Unification: OmniGen not only demonstrates text-to-image generation capabilities but also inherently supports various downstream tasks, such as image editing, subject-driven generation, and visual-conditional generation. 2) Simplicity: The architecture of OmniGen is highly simplified, eliminating the need for additional plugins. Moreover, compared to existing diffusion models, it is more user-friendly and can complete complex tasks end-to-end through instructions without the need for extra intermediate steps, greatly simplifying the image generation workflow. 3) Knowledge Transfer: Benefit from learning in a unified format, OmniGen effectively transfers knowledge across different tasks, manages unseen tasks and domains, and exhibits novel capabilities. We also explore the model’s reasoning capabilities and potential applications of the chain-of-thought mechanism. This work represents the first attempt at a general-purpose image generation model, and we will release our resources at https://github.com/VectorSpaceLab/OmniGen to foster future advancements.
1 Introduction
Figure 1: OmniGen can flexibly follow instructions to complete various tasks, without the need for any additional plugins. For complex tasks (e.g., examples in the second line), it can also be completed end-to-end without cumbersome intermediate steps.
The pursuit of Artificial General Intelligence (AGI) has intensified the demand for generative foundation models capable of handling various tasks within a single framework. In the field of Natural Language Processing (NLP), Large Language Models (LLMs) have become exemplary in achieving this goal, demonstrating remarkable versatility across numerous language tasks. However, the realm of visual generation has yet to reveal a counterpart that matches the universality of LLMs. Current image generation models have demonstrated proficiency in specialized tasks. For instance, in the text-to-image generation filed, models such as the SD series [55, 51, 12], DALL-E [54], and Imagen [25] have made significant strides. Meanwhile, lots of efforts have been made to extend the capabilities of diffusion models for specific tasks, such as ControlNet [75], InstandID [64], and InstructPix2Pix [3]. Currently, for each new task, designing a specific module and fine-tuning it is necessary, which hinders the development of image generation. What’s worse, these task-specific models lead to a significant issue: we cannot accomplish tasks simply through input instructions but require a complex and cumbersome workflow involving multiple models and numerous steps. For example, ControlNet needs to use a detector to extract conditions (such as pose estimation maps) and then loads the corresponding condition module. InstantID can only process single-person images and requires a face detection model to detect faces beforehand, followed by using a face encoder to encode the face. If users want to edit the generated images, additional models like InstructPix2Pix need to be loaded.
Figure 2: The framework of OmniGen. Texts are tokenized into tokens, while input images are transformed into embedding via VAE. OmniGen can accept free-form multi-modal prompts and generate images through the rectified flow approach.
Can a single model complete various tasks end-to-end through user instructions, similar to how ChatGPT handles language tasks? We envision a future where image generation is made simple and flexible: that any tasks can be accomplished directly through user instructions. Motivated by this goal, we propose a unified framework: OmniGen. As shown in Figure 1, this framework allows for convenient and flexible image generation for any purposes, where no additional plugins or operations are needed. Given the flexibility in following arbitrary instructions, the new framework also help to inspire more interesting image generation tasks.
Unlike popular diffusion models, OmniGen features a very concise structure, comprising only two main components: a VAE and a transformer model. OmniGen supports arbitrarily interleaved text and image inputs as conditions to guide image generation, instead of only accepting pure text or image. To train this architecture as a unified model, we construct a large-scale multi-task image generation dataset X2I, which unifies different tasks with one uniform format. We evaluate the well-trained model based on multiple benchmarks, demonstrating its superior generation capability compared to existing models. Remarkably, OmniGen enables effective knowledge transfer across different scenarios, allowing it to handle unseen tasks and domains while also fostering the emergence of new abilities.
Our contributions are summarized as follows:
  • We introduce OmniGen, a unified image generation model that excels in multiple domains. The model demonstrates competitive text-to-image generation capabilities and inherently supports a variety of downstream tasks, such as controllable image generation and classic computer vision tasks. Furthermore, it can handle complex tasks end-to-end without any lengthy intermediate steps. To our knowledge, OmniGen is the first image generation model to achieve such comprehensive functionality.
  • We construct the first comprehensive image generation dataset named X2I, which stands for ”anything to image”. This dataset covers a wide range of image generation tasks, all of which are standardized into a unified format.
  • By unified training on the multi-task dataset, OmniGen can apply learned knowledge to tackle unseen tasks and domains, as well as exhibit new capabilities. Additionally, we explore the model’s reasoning capabilities and chain-of-thought mechanism.
  • OmniGen takes an initial step toward a foundational model for general image generation. We will open-source the relevant resources (model, code, and data) and hope this work can provide some insights for future image generation models.
2 OmniGen
In this section, we present the details of OmniGen framework, including the model architecture and training method.
2.1 Model Design
Network Architecture. The design principles of OmniGen are universality and conciseness. As illustrated in Figure 2, OmniGen adopts an architecture comprised of a Variational Autoencoder (VAE) [27] and a pre-trained large transformer model. Specifically, we use the VAE from SDXL [51] to extract continuous visual features from images. The transformer model initialized by Phi-3 [1] generates images based on instructions that serve as conditions. Only VAE is frozen during training. Unlike state-of-the-art diffusion models that require additional encoders to pre-process conditional information (such as CLIP text encoder and image encoder), OmniGen inherently encodes conditional inputs by itself, significantly simplifying the pipeline. Furthermore, OmniGen jointly models text and images within a single model, rather than independently modeling different input conditions with separate encoders as in existing works [68, 70, 72, 64, 8] which lacks interaction between different modality conditions.
Input Representation. The input to the model can be multi-modal interleaved text and images in free form. We utilize the tokenizer of Phi-3 to process text. For images, we firstly employ a VAE with a simple linear layer to extract latent representations. Then, they are flattened into a sequence of visual tokens by linearly embedding each patch in the latent space. Following [49], the patch size is set to 2. We apply standard frequency-based positional embeddings to visual tokens, and process images with varying aspect ratios as SD3 [12]. In addition, we encapsulate each image sequence with two special tokens: “$<$img$>$” and “$<$/img$>$” before inserting it into the sequence of text tokens. We also append the timestep embedding [49] at the end of the input sequence. We do not need any task-specific special tokens to indicate the task type.
Attention Mechanism. Different from text, which can be decomposed into discrete tokens to model, we argue that the image should be modeled as a whole based on its nature. Therefore, we modify the common causal attention mechanism in LLM by integrating it with the bidirectional attention as illustrated in Figure 2. Specifically, we apply causal attention to each element in the sequence, but apply bidirectional attention within each image sequence. This allows each patch to pay attention to other patches within the same image, while ensuring that each image can only attend to other images or text sequences that have appeared previously.
Inference. During inference, we randomly sample a Gaussian noise and then apply the flow matching method to predict the target velocity, iterating multiple steps to obtain the final latent representation. Finally, we use a VAE decoder to decode the latent representation into the predicted image. Thanks to the attention mechanism, OmniGen can accelerate inference like LLMs by using kv-cache: storing previous and current key and value states of the input conditions on the GPU to compute attention without redundant computations.
2.2 Training Strategy
Train objective. In this work, we use rectified flow [39] to optimize the parameters of model. Different from the DDPM [24], flow matching conducts the forward process by linearly interpolating between noise and data in a straight line. At the step $t$, $\mathbf{x}_{t}$ is defined as: $\mathbf{x}_{t}=t\mathbf{x}+(1-t)\boldsymbol{\epsilon}$, where $\mathbf{x}$ is the original data, and $\boldsymbol{\epsilon}\sim\mathcal{N}(0,1)$ is the Gaussian noise. The model is trained to directly regress the target velocity given the noised data $\mathbf{x}_{t}$, timestep $t$, and condition information $c$. Specifically, the objective is to minimize the mean squared error loss:
$\mathcal{L}=\mathbb{E}\big{[}||(\mathbf{x}-\boldsymbol{\epsilon})-v_{\theta}\big{(}\mathbf{x}_{t},~{}t,~{}c\big{)}||^{2}\big{]}.$ (1)
For image editing tasks, the objective is to modify specific regions of the input image while keeping other areas unchanged. Therefore, the difference between the input image and the target image is often small, which allows the model to learn an unexpected shortcut: simply copying the input image as the output to make the related training loss very low. To mitigate this phenomenon, we amplify the loss in the regions of the image where changes occur. More specifically, we calculate the loss weights for each region based on latent representations of input image $\mathbf{x^{\prime}}$ and target image $\mathbf{x}$:
$\quad w_{i,j}=\begin{cases}1&\text{if }\mathbf{x}_{i,j}=\mathbf{x^{\prime}}_{i,j}\\ \frac{1}{||\mathbf{x}-\mathbf{x^{\prime}}||^{2}}&\text{if }\mathbf{x}_{i,j}\neq\mathbf{x^{\prime}}_{i,j}\end{cases}$ (2)
Consequently, regions with alterations are assigned significantly higher weights than those without changes, guiding the model to focus on the areas to be modified.
Training Pipeline. Following previous work [12, 16, 5], we gradually increase the image resolution during the training process. Low resolution is data-efficient, while high resolution can enhance the aesthetic quality of the generated images. Detailed information for each training stage is presented in Table 1. We adopt the AdamW [40] with $\beta=(0.9,0.999)$ as the optimizer. All experiments are conducted on 104 A800 GPUs.
Stage Resolution Steps (K) Batch Size Learning Rate
1 256$\times$256 500 1040 1e-4
2 512$\times$512 300 520 1e-4
3 1024$\times$1024 100 208 4e-5
4 2240$\times$2240 30 104 2e-5
5 Multiple 80 104 2e-5
Table 1: Details about each training stage of OmniGen.
3 X2I Dataset
To achieve robust multi-task processing capabilities, it is essential to train the model on large-scale and diverse datasets. However, in the field of image generation, a readily available large-scale and diverse dataset has yet to emerge. In this work, we have constructed a large-scale unified image generation dataset for the first time, which we refer to as the X2I dataset, meaning ”anything to image”. We have converted all data into a unified format, and Figure 3 presents some examples of the X2I dataset. The entire dataset comprises approximately 0.1 billion images. A detailed description of the composition will be provided in the following sections.
Figure 3: Examples of X2I dataset. We standardized the input of all tasks into an interleaved image-text sequence format. The placeholder ||||image_i|||| represents the position of the $i$-th input image.
3.1 Text to Image
The input for this subset of data is plain text. We collect multiple open-source datasets from various sources: Recap-DataComp [33](a subset of 56M images), SAM-LLaVA [5], ShareGPT4V [6], LAION-Aesthetic [57](a subset of 4M images), ALLaVA-4V [4], DOCCI [45], DenseFusion [34] and JourneyDB [59]. While these datasets are large in quantity, their image quality is not always high enough. In the early stages of training, we use them to learn a broad range of image-text matching relationships and diverse knowledge. After stage 3, we utilize an internal collection of 16M high-quality images to enhance the aesthetic quality of the generated images. We use the InternVL2 [10] to create synthetic annotations for internal data and LAION-Aesthetic.
3.2 Multi-modal to Image
Different from existing diffusion models, OmniGen can accept more flexible multi-modal instruction, thus adapting to a variety of data and task types.
3.2.1 Common Mixed-modal Prompts
The input of this portion of data is arbitrarily interleaved text and images. We collect the data from multiple tasks and sources: image editing (MagicBrush [74], InstructPix2Pix [3] and SEED-edit [17], ), human motion (Something-Something [21]), virtual try-on (HR-VITON [30] and FashionTryon [77]), and style transfer (StyleBooth [23]). The issue of utilizing additional visual conditions for finer-grained spatial control has garnered widespread attention[75, 31]. We employ the MultiGen [52] dataset to learn this function, and select six representative visual conditions: Canny, HED, Depth, Skeleton, Bounding Box, and Segmentation. These types of tasks take text prompts with specific visual conditions (such as segmentation maps, and human pose maps) as multi-modal inputs. We standardize all tasks into the input-output pair format as shown in Figure 3-(b).
3.2.2 Subject-driven Image Generation
We construct both a large-scale foundational dataset (GRIT-Entity dataset) and a high-quality advanced dataset (Web Images dataset) for subject-driven image generation. For the GRIT-Entity dataset, we leverage the GRIT dataset [50], which annotates object names within images. Using these annotations, we apply GroundingDINO [38] for open-set text-to-bounding-box detection. Based on the grounded bounding boxes, we employ SAM [28] to segment object masks from cropped images. We further use the MS-Diffusion [66] to repaint the object images for higher quality and diversity of data. The process of data construction and the final instruction format are illustrated in supplementary material. Through these methods, we acquire 6M data pairs.
Although the GRIT-based approach provides a substantial amount of samples, the input data extracted directly from original images will tend to lead the model to fall into a simple copy-and-paste pattern. To fully unleash the subject-driven image generation capability of OmniGen, we construct a high-quality natural images dataset from the web. We use the search engine to collect multiple different images of the same person, using one as the input and another as the output. An example is shown in Figure 3-(c). More details about the construction process are given in the supplementary material.
3.2.3 Computer Vision Tasks
We introduce classic computer vision tasks to expand the boundaries of the model’s generative capabilities. For low-level vision tasks (low-light image enhancement [67], deraining [73], deblurring [44], inpainting [52], outpainting [52] and colorization [57]), where the annotation itself is an image, we only add text instructions, which are randomly sampled from instructions generated by GPT-4o. For high-level vision tasks, we choose to represent all annotations as images. We use images from LAION [57] as source images and annotations from [52] as targets to construct image pairs (such as source image and its human pose mapping). The annotations include human pose, depth mapping, canny, Hed edge, and segmentation. Additionally, we also use several datasets for referring image segmentation, including RefCOCO [26], ADE20k [78], and ReasonSeg [29]. As shown in Figure 3-(d), the input is the source image and a natural language expression, the output is an image with the corresponding object highlighted in blue.
3.3 Few-shot to Image
We build a few-shot to image dataset to stimulate the in-context learning ability of the model. Specifically, for each task described in the preceding sections, we randomly selected a few examples and combined the original input with these examples to form new inputs. The specific data format is shown in Figure 3-(e). Due to limitations in training resources, we opted to use only one example to improve training efficiency.
4 Experimental Results
The default inference step is set to 50. We use the classifier-free guidance for both text condition and image condition following  [3]. The default guidance scale is set to 2.5, and the image guidance scale is set to 1.6. For image editing task, we increase the image guidance scale to 2.
4.1 Image Generation
4.1.1 Qualitative Results
Figure 4: The results of different image generation tasks.
Figure 5: Subject-driven generation. When the reference image contains multiple objects, OmniGen can automatically identify the required objects based on textual instructions without the need for additional preprocessing steps like image cropping or face recognition.
Figure 4 summarizes the results of different image generation tasks, including text-to-image, image editing, and visual conditional. These results demonstrate that OmniGen can handle various downstream tasks based on multi-modal instructions.
Figure 5 presents the outcomes of the subject-driven generation task. OmniGen can extract the required objects from the given reference images and generate new images accordingly. Furthermore, when the reference image contains multiple objects, the model can directly select the needed objects based on textual instructions (e.g., the one wearing the dark red shirt in image) without requiring additional preprocessing steps. In contrast, existing models [64, 22] require manually cropping the target person from a multi-person image, detecting faces with a face detector, encoding faces with a face encoder, and inputting them into the diffusion model through a cross-attention module. Without these complex processes, OmniGen can complete the task in a single end-to-end step.
More qualitative results are provided in supplementary materials.
4.1.2 Text to Image
Following [12], we evaluate text-to-image generation capability on the GenEval [20] benchmark. We compare the performance of OmniGen with the reported results of other popular image generation models, as summarized in Table 2. Surprisingly, OmniGen achieves similar performance compared to the current diffusion models, such as SD3, which underscores the effectiveness of our framework. Notably, OmniGen has only 3.8 billion parameters, whereas the SD3 model has a total of 12.7 billion parameters. The architecture of OmniGen is significantly simplified, eliminating the cost of an additional encoder, thereby greatly enhancing the efficiency of parameter utilization.
Single Two Attribute
Model Params Overall object object Counting Colors Position binding
SDv1.5 0.9B+0.1B 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SDv2.1 0.9B+0.4B 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL 2.6B+0.8B 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 3.5B+1.0B 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 0.67 0.96 0.87 0.47 0.83 0.43 0.45
IF-XL 5.5B+4.7B 0.61 0.97 0.74 0.66 0.81 0.13 0.35
SD3 8.0B+4.7B 0.68 0.98 0.84 0.66 0.74 0.40 0.43
OmniGen 3.8B 0.70 0.99 0.86 0.64 0.85 0.31 0.55
Table 2: Results of GenEval Benchmark. means the parameter of the frozen text encoder. Our model achieves comparable performance with a relatively small parameter scale.
4.1.3 Multi-modal to Image
[…]
I-Pix2Pix [3] 0.219 0.834
MagicBrush [74] 0.222 0.838
PnP [63] 0.089 0.521
Null-Text [42] 0.236 0.761
EMU-Edit [58] 0.231 0.859
OmniGen 0.231 0.829
[…]
Tex-Inv [13] 0.255 0.780
DreamBooth [56] 0.305 0.803
ReImagen [7] 0.270 0.740
SuTI [9] 0.304 0.819
Kosmos-G [47] 0.287 0.847
OmniGen 0.315 0.801
Table 3: Left: Results on EMU-Edit test data. Right: Results on DreamBench. As a universal model, OmniGen demonstrates performance comparable to that of the best proprietary models.
Seg. Mask Canny Edge Hed Edge Depth Map
[…]
T2I-Adapter [43] 12.61 23.65 - 48.40
Gligen [35] 23.78 26.94 0.5634 38.83
Uni-ControlNet [76] 19.39 27.32 0.6910 40.65
UniControl [52] 25.44 30.82 0.7969 39.18
ControlNet [75] 32.55 34.65 0.7621 35.90
ControlNet++ [31] 43.64 37.04 0.8097 28.32
OmniGen 40.06 38.96 0.8332 31.71
Table 4: Comparison with SOTA methods on controllable image generation. $\uparrow$ indicates higher result is better, while $\downarrow$ means lower is better.
We evaluate the image editing on EMU-Edit [58] dataset and subject-driven generation capability on DreamBench [56]. We use CLIP-T to measure how well the model followed the instructions, while CLIP-I similarity scores measure the model’s ability to preserve elements from the source image. As shown in Table 3, OmniGen exhibits comparable performance to the current state-of-the-art models. Noted for subject-driven generation task, we select one image of the specific object as input instead of fine-tuning the model like DreamBooth.
In Table 4, we use the dataset and script from  [31] to evaluate the generation capability based on visual conditions. For each condition, the controllability is evaluated by measuring the similarity between the input conditions and the extracted conditions from generated images of diffusion models. We can find that OmniGen achieves competitive results for all conditions.
Figure 6: The results of OmniGen in various traditional vision tasks.
Figure 7: ControlNet involves multiple steps and plugins: firstly using the corresponding detector to extract information from the reference image, and then loading the appropriate ControlNet module to process the visual conditions. In contrast, OmniGen completes the entire task in a single step within a single model.
4.2 Computer Vision Tasks
We present several qualitative results of computer vision tasks in Figure 6. OmniGen can handle various low-level vision tasks such as deraining, deblurring, and inpainting. At the bottom of Figure 6, we can see that OmniGen is also able to handle high-level vision tasks, such as human pose recognition.
Note that OmniGen’s capabilities in CV tasks are not intended to surpass existing state-of-the-art models that developed for specific tasks over a long period. We hope it can transfer the knowledge gained from these traditional computer vision tasks to image generation tasks, thereby unlocking greater potential. Some examples are shown in Figure 7. The existing workflow for ControlNet involves using a detector to extract spatial condition information from the reference image, and then loading the corresponding control module to model the spatial condition information for image generation. Can we directly use OmniGen to generate new images based on a reference image in only one step? Surprisingly, even without having encountered such a task before, OmniGen handles it admirably without explicit extraction of additional conditional images. More specifically, we can directly input the reference image and text instruction (e.g., Follow the human pose of this image to generate new image.) to generate the target image in only one step without any explicit additional intermediate procedures.
5 Further Analysis
LLMs demonstrate remarkable generalization capabilities, and can boost performance through mechanisms such as in-context learning and chain of thought. We observe similar functionalities in OmniGen as well, and present our findings in this section.
5.1 Emgerging Capabilities
By standardizing all tasks into a unified format and training on X2I dataset, OmniGen can acquire universal knowledge and allow knowledge transfer across different scenarios and tasks, thus enabling the generation capabilities on unseen tasks and domains.
Figure 8: Examples of Emgerging Capabilities of OmniGen.
Task Composition. In real-world applications, user requirements often involve combinations of tasks. As shown in Figure 8-(a), our model is capable of simultaneously processing multiple instructions, including those for different tasks as well as multiple instructions for the same task. These results highlight our model’s versatility and potential for widespread adoption in the wild.
End-to-end Workflow. Users typically need to load multiple models and perform multi-step processing to ultimately generate a satisfactory image, making the workflow very cumbersome and costly. OmniGen possesses both excellent multi-modal understanding and image generation capabilities, enabling it to complete a lot of tasks without relying on external models, thereby significantly simplifying the workflow and saving cost. As demonstrated in Figure 7, OmniGen can extract the relevant conditional information from the reference image and generate a new image based on the captured condition within one step. This process is implicit, with all processing completed internally within the model, only requiring the user to input a simple command.
In-context Learning for Unseen Domains. We use the data from FSS [32] which contains objects that have never been seen in previous datasets to evaluate the generalized ability. In the right of Figure 8-(b), we can see that OmniGen is not familiar with the concepts of “pencil sharpeners”, and it cannot identify it from images. However, when provided with an example, the model is capable of making accurate predictions. By providing an example, the model is capable of successfully completing this unseen segmentation task.
5.2 Reasoning Ability
We have explored the reasoning capabilities of the model and presented the results in Figure 9. As shown in the left half of Figure 9, when given an instruction without explicitly specifying the object, such as “Where can I wash my hands? Please help me find the right place in image_1”, the model can recognize image contents and infer that a sink is needed. Consequently, the model identifies and indicates the area of the sink in the image. This functionality creates potential applications in the field of embodied intelligence, assisting intelligent agents in comprehending multi-modal instructions, locating necessary objects and planning subsequent actions. Moreover, the right half of Figure 9 demonstrates that after inferring the target object, the model can also perform editing operations on it. If no object matches, the model will refrain from editing any unrelated objects(see the example in the bottom right corner of Figure 9).
Figure 9: Reasoning ability of OmniGen.
5.3 Chain of Thought
Figure 10: Simulate the process of human drawing via step-by-step generation.
The Chain-of-Thought (CoT) method can significantly boost the performance of LLMs by decomposing the task into multiple steps and sequentially solving each step to obtain an accurate final answer. We consider whether a similar alternative can be applied to image generation. Inspired by the basic way of human drawing, we hope to mimic the step-by-step drawing process, iteratively refining the image from a blank canvas. We construct an anime image dataset and use the PAINTS-UNDO [^1] https://github.com/lllyasviel/Paints-UNDO model to simulate each stage of artwork creation. We select 8 representative frames to depict the gradual development of the final image. After filtering out inconsistent sequences, we fine-tuned the model on this dataset for 16,000 steps.
The results are visualized in Figure 10, alongside the outputs generated by the original model. For step-by-step generation, the input data consists of the current step’s image and text, and then the model predicts the image for the next step. It can be observed that the fine-tuned model successfully simulates the behavior of a human artist: drawing the basic outline, incrementally adding details, making careful modifications, and applying colors to the image. In this manner, users can modify the previous results to control the current output, thereby participating more actively in the image generation process, rather than passively waiting for the final image with a black-box diffusion model. Unfortunately, the quality of the final generated images does not surpass that of the original model. In the step-by-step generation approach, the model may incorporate erroneous modifications, leading to some disarray in the final image. This does not imply that the approach is unfeasible; currently, we only conduct a preliminary exploration, leaving further optimizations for future research. Based on the findings of previous work [36] on LLMs, which indicate that process supervision significantly outperforms outcome supervision, we posit that supervising the drawing process of images is a promising direction that may assist the model in handling more complex and diverse scenes.
6 Ablation Study
In this section, we assess the necessity of certain modules in OmniGen
Attention Mask. Figure 11-Top shows the generated images without the modified attention in Section 2.1. As we can see, there is a lot of noise in the images. Compared to Figure 4, the generated images in Figure 11-Top have many distorted areas, confirming that the unidirectional attention mechanism in LLMs is not suitable for the rectified flow.
Weighted Loss. Figure 11-Bottom shows the results when the weighted loss from Section 2.2 is not used for training. For image editing tasks, since the edited area is usually small, the model tends to learn to copy the input image directly as the output. Comparing Figure 11-Bottom and Figure 4, it’s evident that using weighted loss can prevent the model from learning this shortcut.
Input Image Representation. The CLIP model is widely used in multi-modal understanding models [37]. We also explored the difference between using CLIP [^2] https://huggingface.co/openai/clip-vit-large-patch14-336 and using VAE to process input images. The results are shown in Table 5. There is not much difference between the two strategies. Using VAE has a slight advantage in image similarity, which might be related to the fact that the output images also require VAE processing. Another advantage of using VAE is that it avoids additional modules, maintaining the simplicity of the model. Directly inputting the original image is another simpler approach that we leave for future exploration.
Figure 11: Top: results of causal attention; Bottom: results without loss weights.
Input Image EMU-Edit DreamBench
Representation CLIP-T CLIP-I CLIP-T CLIP-I
CLIP 0.232 0.820 0.312 0.789
VAE 0.231 0.829 0.315 0.801
Table 5: Comparison of different methods to obtain visual embedding for input images.
7 Related Work
Generative Foundation Models. The GPT series [53, 46] have demonstrated that language models can learn numerous tasks via training on a large-scale dataset. Beyond language, multi-modal large language models [37, 11] have been proposed to integrate vision and language capabilities. However, they lack the capability to generate images. Some works propose integrating LLMs with diffusion models to equip LLMs with image generation capability [60, 18, 69, 61]. Others use discrete tokens to support both image and text generation simultaneously [62, 41, 71]. They focus on multi-modal generation with limited image-generation capability. Concurrent works such as TransFusion [79] and Show-O [71] unify diffusion and autoregressive methods into a single model, generating text autoregressively and images through diffusion. Nonetheless, like existing diffusion models, these works focus on a limited range of image generation tasks, primarily text-to-image generation, and cannot cover more complex and various visual generation tasks. The construction of a universal foundation model for image generation remains unclear and has not been fully explored.
Diffusion Model. Recent advancements in diffusion models have been remarkable, with notable contributions from the Stable Diffusion (SD) series [55, 51, 12], DALL-E [54], and Imagen [25]. These models are predominantly designed for text-to-image generation tasks. Many efforts have been made to extend the capabilities of diffusion models, such as ControlNet [75], T2I-Adapter [43], StyleShot [15]. InstructPix2Pix [3], InstructSD [48] and EMU-edit [58] explore performing general image editing tasks through instructions. However, these methods are task-specific, extending the capabilities of SD by modifying the model architecture. In contrast, OmniGen is a model that natively supports various image-generation tasks, and no longer requires any preprocessing steps or assistance from other models. There is some work exploring the unification of computer vision (CV) tasks [2, 65, 14, 19]. However, these efforts primarily focus on classic vision tasks instead of general image generation tasks, and often underperform compared to those specifically designed and trained for corresponding tasks. In our work, the introduction of CV tasks does not aim to obtain a SOTA performance on CV tasks, which is designed to enable the model to learn general knowledge.
8 Limitations and Conclusion
In this work, we introduce OmniGen, the first unified image generation model. It is designed to be simple, flexible, and easy to use. We also construct the first large-scale unified image generation dataset, X2I, to activate the general capabilities of the OmniGen model. OmniGen has demonstrated excellent image-generation abilities across various tasks.
However, the current model still has some issues. The text rendering capability is limited, and long text can not be accurately generated. The output images may contain undesired details (e.g. abnormal hand). Besides, previously unseen image types (e.g., surface normal map) can hardly be processed as expected. More failure cases are provided in supplementary materials.
We believe that the future paradigm of image generation should be simple and flexible, allowing for the direct generation of various images through any multimodal instructions without complex workflows. OmniGen represents an important step towards a foundational model for universal image generation. We will open-source the relevant resources, hoping to provide insights for the future of image generation.
Supplementary Material
9 More Qualitative Results
In this section, we provide more qualitative results of OmniGen. Overall, OmniGen achieves competitive or even superior results compared to task-specific open-source models. It’s important to note that OmniGen is a general-purpose model and is not specifically trained for any particular task. Additionally, OmniGen has capabilities that many other models lack, allowing it to follow multimodal instructions to complete complex tasks. To the best of our knowledge, OmniGen is the first model capable of such flexible control in image generation.
9.1 Text to Image
Figure 12 shows more examples of text-to-image task. OmniGen can generate images with arbitrary aspect ratios.
9.2 Image Editing
In Figure 13, we provide qualitative results of OmniGen compared to some other open-source models. As shown, OmniGen achieves better results than existing open-source models, accurately modifying specified targets without altering other content. Emu-edit is not open-source, so further evaluation was not possible. Additionally, OmniGen demonstrates strong instruction comprehension capabilities, handling complex tasks such as executing two editing instructions in one step, as seen in Figure 13-(e), and reasoning editing tasks in the (f). In Figure 13-(f), we don’t not explicitly specify the target to be modified; the model infers the target (balloon) from the description and makes precise modifications. OmniGen inputs images and text into a unified transformer, allowing for extensive interaction between different modalities, resulting in powerful multimodal instruction following capabilities, which other frameworks like InstructPix2Pix cannot achieve.
9.3 Subject-driven
Figure 14 shows more examples of subject-driven generation on DreamBench. Similar to Kosmos-G, we use an image of an object as a reference and generate a new image based on the text description, without fine-tuning on the specific object. As illustrated, OmniGen achieves better subject fidelity and text fidelity, meaning it preserves the object from the reference image while following new textual instructions. Kosmos-G struggles to maintain the object from the reference image. Notably, Kosmos-G does not always adhere to the text instructions; as shown in Figure 14-(d), it fails to depict the described content. Instead, Kosmos-G nearly replicates the reference image, which is why it has a higher CLIP-I similarity (see Table 3 in the main paper). Additionally, Kosmos-G uses SD1.5 as the image model, resulting in lower image quality compared to OmniGen.
9.4 Identity-Preserving Generation
Figure 15 shows the results of Identity-Preserving generation task. OmniGen achieves better quality than instandID and comparable results to the latest model PULID-FLUX  [^3] https://huggingface.co/spaces/yanze/PuLID-FLUX. It should be noted that as a general model, OmniGen’s facial similarity is not yet on par with the latest specialized model, PULID-FLUX, which is an area for future improvement. Unlike these models, which require a face detector to detect faces and a face encoder to extract facial features, OmniGen does not need task preprocessing steps or additional plugins. It automatically completes tasks with input images and instructions. Moreover, OmniGen does not only focus on faces; it can capture other useful reference information in the input image, such as clothing. For example, in Figure 15-(a) and -(b), OmniGen tends to preserve the clothing of the person in the input image. Of course, when clothing is specified in the instructions, OmniGen can also follow the instructions to use new clothing (see Figure 15-(c) and -(d)). In contrast, other models can only use the extracted face and cannot utilize the original clothing from the input image.
Figure 16 showcases some unique features of OmniGen. Existing models cannot handle input images with multiple objects or multiple input images. OmniGen can extract desired individuals from a group photo based on instructions, and it can also process multiple images simultaneously, as shown in the Figure 16-(a). Besides, OmniGen can handle more than just people; as demonstrated in Figure 16-(b), it can identify specific clothing items in an image and use them to generate new images.
9.5 Failure cases
Figure 17 illustrates several typical failure cases of the current model. We summarize the limitations of the current model as follows:
  • Similar to existing diffusion models, OmniGen is sensitive to text prompts. Typically, detailed text descriptions result in higher-quality images.
  • The current model’s text rendering capabilities are limited; it can handle short text segments but fails to accurately generate longer texts. Additionally, due to resource constraints, the number of input images during training is limited to a maximum of three, preventing the model from handling long image sequences.
  • The generated images may contain erroneous details, especially small and delicate parts. In subject-driven generation tasks, facial features occasionally do not fully align. OmniGen also sometimes generates incorrect depictions of hands.
  • OmniGen cannot process unseen image types (e.g., image for surface normal estimation).
We believe that most limitations can be addressed by training the model on more related data. Moreover, compared to most models, fine-tuning OmniGen for downstream tasks is simpler. Due to OmniGen’s inherent support for multimodal inputs, users do not need to spend significant effort on designing networks for specific tasks (for example, there is no need to add ControlNet for conditional generation, nor to design additional style encoders for image styles). We will open-source the training code, and users will only need to prepare their data to fine-tune it for various tasks, including new tasks that have never been supported by any model before. We believe this approach can inspire a wide range of interesting downstream applications.
Due to a lack of high-quality image data and insufficient model size, OmniGen’s image quality isn’t as good as the latest FLUX model. On one hand, expanding the data and model size is crucial; on the other hand, enabling existing models like FLUX to have similar multimodal instruction-following capabilities is also a potential research direction.
Figure 12: Examples for text-to-image task. OmniGen can generate images with arbitrary aspect ratios.
Figure 13: Qualitative results of image editing task. OmniGen achieves better results than other open-source models. Thanks to modeling both text and images within a unified transformer, OmniGen demonstrates strong multimodal instruction understanding and is capable of handling complex tasks, as shown in (f).
Figure 14: Qualitative results on DreamBench. OmniGen achieves better subject fidelity and text fidelity, preserving the object from the reference image while following new textual instructions. In contrast, Kosmos-G does not always follow text instructions, resulting in a higher CLIP-I similarity.
Figure 15: Results of Identity-Preserving Generation. OmniGen doesn’t require additional face detectors and encoders. OmniGen can utilize clothing information from input images, as shown in examples (a) and (b), and can also follow text instructions to generate new clothing, as seen in examples (c) and (d), demonstrating a higher degree of flexible control.
Figure 16: Some unique features of OmniGen. OmniGen can not only extract the desired object from image containing multiple objects based on instructions, but also process multiple images simultaneously. To our knowledge, no other model currently can offer this level of flexible and precise control.
Figure 17: Failure cases of OmniGen
Figure 18: Left: Illustration of the construction process for the GRIT-Entity dataset. We used instance segmentation and repainting methods to acquire a large volume of data. Right: Illustration of the cross-verification strategy used in constructing our web images dataset. For a group photo of Person-A and Person-B, we sampled several images from individual photos of Person-A and Person-B and asked MLLM whether they appear in the group photo. A group photo is retained only if the ”Yes” ratio for both Person-A and Person-B meets a specific threshold. The individual images marked as ”Yes” are then used to construct data pairs with the corresponding group image.
10 Subject-driven dataset
Figure 18-Left shows the process of constructing the GRIT-Entity dataset. Although the GRIT-based approach provides a substantial amount of data, the input data extracted directly from original images can lead the model to fall into simple copy-paste patterns. To fully unleash the subject-driven image generation capability of OmniGen, we constructed a high-quality web images training dataset using natural images of well-known individuals. First, we sampled 20 million Alt-text entries from the Datacomp dataset [^4] https://huggingface.co/datasets/UCSC-VLAA/Recap-DataComp-1B and used spaCy [^5] https://github.com/explosion/spaCy for named entity recognition. We selected the most frequently occurring names and employed GPT-4o to filter out real, notable individuals, resulting in 2,000 names. Furthermore, we expanded the initial 2,000 names by including closely related individuals, resulting in approximately 10,000 name pairs. We then scraped images of these individuals and pairs from search engines. Due to the noise in web images, where scraped images may not contain the specified individuals, we designed a cross-verification strategy using InternVL [10] to filter single and group images, as detailed in Figure 18-Right. The retained single and group images were then captioned with details such as attire and actions. We successfully build a dataset of 533K image pairs.